Equipment acquired at a cost of Rs. 72,000 has an estimated residual value of Rs. 3,000 and an estimated useful life of 5 years. It was placed in service on July 31 of the current fiscal year, which ends on December 31, Determine the depreciation for the current fiscal and the following fiscal year by

(a) by the straight line method,

(b) by the declining-balance method, using twice the straight line rate (40%) and

(c) by the sum of the year digit method.


1
Expert's answer
2021-04-26T19:34:07-0400

a)


"\frac{5}{12}\times\frac{72000-3000}{5}=5750"

b)


"\frac{5}{12}\times 0.4\times (72000-3000)=11500"

c)


"\frac{5}{12}\times \frac{72000}{15}=2000"